A uniform 5.0 m long ladder of mass 25 kg leans against a frictionless wall at an angle of 62 degrees with the floor. Determine the minimum coefficient of static friction needed on the floor for the ladder not to slide.
1
Expert's answer
2019-11-28T08:35:44-0500
"N_F=mg, N_W=F_{friction}=\mu N_F=\mu mg"

"N_wL\sin{\phi}=mg(0.5L)\cos{\phi}"

So,


"\mu mgL\sin{\phi}=mg(0.5L)\cos{\phi}"

"\tan{\phi}=\frac{1}{2\mu}"

"\tan{62}=\frac{1}{2\mu}"

"\mu=0.27"